OpenStudy (anonymous):

Solve. 10/(3x) + 4/3 = (7 + x)/2x A) x = 1/3 B) x = 17/5x C) x = 1/5 D) x = 1/6 help??

OpenStudy (anonymous):

well first expand the right side

OpenStudy (anonymous):

10/(3x) + 4/3 = (7 + x)/2x multiply through by x 10/3 +4x/3 =(7+x)/2 multiply through by 6 to get rid of the denominators 10*2 + (4x)(2) = (7+x)(3) expand and solve for x. it should be doable now.

OpenStudy (anonymous):

@tomo it isn't, I'm sorry I'm just so bad at math! :-( can you help me through this entire equation so I can try the next by myself?

OpenStudy (anonymous):

10*2 + (4x)(2) = (7+x)(3) 20 + 8x = 21 + 3x 8x-3x = 21-20 5x = 1 x = ?

OpenStudy (anonymous):

@tomo so it's A??

OpenStudy (anonymous):

please write down what you are doing, because otherwise i can't tell what you are struggling with or thinking.

OpenStudy (anonymous):

@tomo I'm just so confused.. please, can you just show me how you do the entire problem

OpenStudy (anonymous):

10/(3x) + 4/3 = (7 + x)/2x multiply through by 6x to get rid of the denominators. (this is the least common denominator) so 6x[10/(3x)] +(6x)(4/3) = (6x)[(7+x)/(2x)] 2*10 + 2x*(4) = 3*(7+x) 20 +8x = 21 + 3x subtract 20 from each side 8x = 21-20 +3x 8x = 1 + 3x subtract 3x from both sides 8x-3x = 1 5x = 1 x = 1/5